OpenStudy (anonymous):

Can someone walk me through solving this problem so I can do the rest of them? Which graph represents the solution to the given system? y = –2x + 5 and y = 3x – 2

OpenStudy (anonymous):

Plug in y=3x-2 where the y is in the equation y-2x=-5 Combine like terms 3x-2-2x=-5 Add to to 2 to both sides x-2=-5 Plug in x=-3 into the equation y=3x-2 y=3•-3-2 Solve for y, y=-9-2 Le Answer: y=-11 (-3,-11)

OpenStudy (anonymous):

Those aren't any of my choices!

OpenStudy (jdoe0001):

why not just graph them both, and find out

OpenStudy (radar):

Something is awry. We know that y=y so we can write: -2x + 5 = 3x -2 using legal algebra we then get -5x = -7 resulting in x=7/5 then solve for y getting y=11/5

OpenStudy (radar):

Verify: 11/5 = -14/5 + 25/5 11/5 = 11/5 verify other equation: 11/5= 3(7/5) - 2 11/5 = 21/5 - 10/5 11/5=11/5 Verified. Now find the graph that match ((7/5, 11/5)

OpenStudy (radar):

Should be two lines intersecting at solution. One line with a negative slope, one with a positive slope,